OpenStudy (mendicant_bias):

I'm trying to solve an IVP using laplace transforms and must specifically use s-axis and t-axis shifting properties, but am running into issues putting the inverse laplace in the right form. Posted below in a moment.

OpenStudy (mendicant_bias):

\[y''-6y'+13y=0, \ \ \ y(0)=0, \ y'(0)=-3\]

OpenStudy (mendicant_bias):

\[S^2Y(s)-sy(0)-y'(0)-6[sY(s)-y(0)]+13 Y(s)=0\]

OpenStudy (sidsiddhartha):

CORRECT SO FAR

OpenStudy (mendicant_bias):

\[Y(s)[s^2-6s+13]=-3\]

OpenStudy (mendicant_bias):

\[Y(s)=\frac {-3}{s^2-6s+13}\]

OpenStudy (mendicant_bias):

\[L^{-1} \left\{\vphantom{} \frac{-3}{(s-\sqrt{6})^2+7} \right\} = L^{-1} \left\{\vphantom{} \frac{\sqrt{7}}{(s-\sqrt{6})^2+7} \right\}-L^{-1} \left\{\vphantom{} \frac{?}{(s-\sqrt{6})^2+7} \right\}\]

OpenStudy (mendicant_bias):

I guess I'm having trouble figuring out how to non-messily do that in the denominator, putting the inverse laplace in the form necessary to use exponential shifting in combination with the sin(kt) form, to do that.

OpenStudy (sidsiddhartha):

\[\frac{ 1 }{ s^2-6s+13 }=\frac{ 1 }{ s^2-2.s.3+3^2+4 }=\frac{ 1 }{ (s-3)^2+4 }\] try this

OpenStudy (sidsiddhartha):

-3 is just a constant term

OpenStudy (sidsiddhartha):

getting this? i will become\[f(t)=-3.L^{-1}\frac{ 1 }{ (s-3)^2+4 }=\frac{ -3 }{ 2 }.L^{-1}\frac{ 2 }{ (s-3)^2+2^2 }\\=\frac{ -3 }{ 2 }.e^{3t}\sin(2t)\] looks good?

OpenStudy (sidsiddhartha):

*it will become :P

OpenStudy (sidsiddhartha):

@Mendicant_Bias

OpenStudy (mendicant_bias):

I have no idea what you did with all those periods in the original post, heh.

OpenStudy (sidsiddhartha):

ok first step: i made (s^2-6s+13)=(s-3)^2+4 ok?

OpenStudy (mendicant_bias):

Oh my god, what the heck was I even thinking!

OpenStudy (mendicant_bias):

Yeah, no, I don't even know where I was going with the root six thing, heh, that makes no sense.

OpenStudy (sidsiddhartha):

yup it was that easy LOL

OpenStudy (mendicant_bias):

Alright, gonna try to finish it now. Man. Need to look out for those errors.

OpenStudy (sidsiddhartha):

yeah :)

OpenStudy (mendicant_bias):

Oh, nevermind, you finished it in your original post, but yeah, I got it. Thank you.

OpenStudy (sidsiddhartha):

at your service :)